OpenStudy (anonymous):

there are 100 gumdrops. 73 of them are large, and the rest are small. 39 of them are red, and the rest are other colors. 19 of the gumdrops are large red ones. you choose 1 gumdrop at random, and it is that yummy red color. what is the probability that the gumdrop is large, given that it is red?

OpenStudy (anonymous):

\[P(\text{large }|\text{ red})=\frac{P(\text{large AND red})}{P(\text{red})}\] You know that \(19\) of the total \(100\) are large and red, and you know that \(39\) of the \(100\) are red.

OpenStudy (anonymous):

yes I figured that much out

OpenStudy (anonymous):

Okay, so the probability is you want the ratio of these proportions: \[P(\text{large }|\text{ red})=\frac{\dfrac{19}{100}}{\dfrac{39}{100}}=\cdots\]

OpenStudy (anonymous):

*you want is*